# Did Translation Models Get More Robust Without Anyone Even Noticing?

### v1

@@ 4 Robustness to Social Media Text
The previous experiments show that large translation models and LLMs are more robust to synthetic character perturbations than conventional MT models. But is this result applicable to “authentically noisy” domains such as social media text? The nature of “noise” here is different than in the synthetic task: social media text does not necessarily contain many errors (Rello and Baeza-Yates, 2012), but the domain is very different from FLORES. These factors make it difficult to isolate the effect of noise from the general domain adaptation problem. In an ideal world, we would have a translation corpus in which each example is a triple consisting of an original noisy source sequence, a manually annotated cleaned source sequence, and a gold standard translation. This would allow translations of clean and noisy versions of the same source to be compared on some reference-based metric, isolating the effect of the errors. Unfortunately, to our knowledge no such corpus exists, so instead we perform two complementary investigations. First, we evaluate our models on MTNT (Michel and Neubig, 2018), a noisy social media translation corpus. Although this is a useful test of our models’ capabilities, the noise in the corpus is not labeled and there is no clean version of the same data to compare to. This motivates our second experiment, in which we translate data from MultiLexNorm (van der Goot et al., 2021), a lexical normalization benchmark. Together, these two experiments allow us to see both which models succeed and how badly they fail.
4.1 MTNT Experiments
MTNT pairs reddit posts with high-quality professional translations. Although the references are somewhat clean, the sources are only lightly filtered, making them potentially noisy. Unfortunately no cleaned sources exist, making the effect of noise difficult to isolate. Despite this difficulty, it is often used as a robustness benchmark (Karpukhin et al., 2019; Park et al., 2020; Salesky et al., 2021; Vaibhav et al., 2019, inter alia).
Connection to synthetic noise.
Michel and Neubig (2018) estimate that percent of English tokens in MTNT enfr are misspelled. This is much less noise than we used in our synthetic experiments, but is higher than in formal corpora.
Finetuning.
We finetuned OPUS and NLLB on the MTNT train sets for enfr, using early stopping with patience 3 and validating by loss every 100 steps. Other hyperparameters are in Table 7.
OPUS | NLLB | |
Initial learning rate | ||
Batch size | ||
Batch type | Tokens | Sequences |
#A6000 GPUs | 1 | 4 |
Max training | 20k steps | 3 epochs |
Method | enfr | fren |
OPUS | 77.51 | 80.07 |
+finetune | 79.37 | 81.67 |
NLLB | 79.20 | 80.70 |
+finetune | 80.85 | 82.60 |
Tower | 80.65 | 82.80 |
ChatGPT | 81.33 | 84.72 |
Results.
Results are shown in Table 8. GPT-3.5 performs best. Among other models, Tower is generally strongest, although NLLB roughly equals it after finetuning. OPUS performs worst.
@@ 4.2 MultiLexNorm Experiments
The MultiLexNorm dataset covers 13 languages, but many of the corpora are tiny or are in languages that Tower does not cover, so we only use English, German, and Spanish151515For experiments involving Spanish, we use Helsinki-NLP/opus-mt-tc-big-en-es (234.8M parameters, 55k vocabulary) and Helsinki-NLP/opus-mt-es-en (77.9M parameters, 65k vocabulary) from transformers. as sources. In experiments with English sources, we translate to German and Spanish; otherwise, we translate to English. Statistics are presented in Table 9.
@@
As MultiLexNorm lacks reference translations, we use three reference-free evaluation techniques. First, we use faux-BLEU (Anastasopoulos, 2019), which computes (Papineni et al., 2002), where is the hypothesis computed from the noisy source, is the hypothesis computed from the clean source, and is treated as a pseudoreference.161616Specifically, we use spBLEU (NLLB Team et al., 2022). By analogy we also compute faux-COMET. These faux-metrics measure the similarity between and , with faux-BLEU being a lexical metric that captures surface-level features, while faux-COMET is more semantic. In addition, we use a new metric that we dub QE. Given , , noisy and clean source sequences and , and a reference-free quality estimation metric QE, . We use COMETKiwi (Rei et al., 2022b) to compute QE. A QE close to zero means that a model produces similar-quality outputs for both inputs, indicating robustness, whereas a large positive value indicates that translation quality suffers on noisy data.
@@ Results.
Table 6 shows the performance of all models with both noisy and gold-standard cleaned versions of the corpora. In terms of QE, GPT-3.5 performs best for all language pairs. It also has the best faux-COMET for all except ende (where Tower passes it). By faux-BLEU it is outperformed by Tower for ende and NLLB for esen. There is a contrast between the performance of NLLB and Tower. While Tower has the better QE for all language pairs and the better faux-COMET for all except esen, NLLB outperforms it by faux-BLEU for all except ende. This suggests that NLLB preserves lexical structure, while Tower preserves “deeper” features.
5 Source Correction
Initial learning rate | |
Batch size | |
Max train steps | k |
Validation steps | |
Patience |
So far we have shown that OPUS is less robust to synthetic noise than larger models and performs worse on social media text. Next we handle noise with source correction pipelines. Pipelines cannot be learned end-to-end and introduce a risk of error propagation, but they are modular and interpretable, and can be used with closed-source models.
5.1 Correction Models
swap | chardrop | chardupe | key | |
@@ 5.1 Correction Models
mT5-Large | 85.8 | 83.2 | 95.0 | 87.9 |
We use two architectures for source correction:
JamSpell (Ozinov, 2019) is a non-neural spell-checker with trigram context. We use the default English model from the package.171717https://github.com/bakwc/JamSpell
We finetune mT5-based (Xue et al., 2021) correctors on the Prob+Word dataset from NeuSpell181818https://github.com/neuspell/neuspell (Jayanthi et al., 2020). We use the mT5-Large (1.2B parameter) version of the architecture with the hyperparameter ranges in Table 10. We select the checkpoint with the best chrF (Popović, 2015) when decoding a noised version of the FLORES English dev set. We decode with a beam size of 5.
These two models are minimalistic and maximalistic approaches to correction. JamSpell is fast (more than 10k tokens per second on a laptop CPU) but uses only trigram context. On the other hand, mT5 is a large encoder-decoder that leverages a long context. Its subword vocabulary allows it to make open-vocabulary corrections, but it is much slower than JamSpell and is prone to hallucination.
5.2 Correcting Synthetic Errors
Model | ende | enfr | enko | enpt |
OPUS | -0.44 | -0.45 | -0.30 | -0.38 |
NLLB | -0.33 | -0.39 | -0.16 | -0.37 |
Tower | -0.28 | -0.30 | 0.00 | -0.26 |
GPT-3.5 | -0.30 | 0.32 | -0.10 | -0.32 |
Model | ende | enfr | enko | enpt |
OPUS | 0.10 | 0.00 | 0.01 | -0.01 |
NLLB | -0.04 | -0.03 | -0.03 | 0.03 |
Tower | -0.04 | -0.08 | -0.03 | -0.07 |
GPT-3.5 | -0.11 | -0.08 | 0.07 | -0.04 |
MT results.
The effect of correcting synthetic enfr errors is shown in Figure 4. For OPUS models, source correction is often beneficial, with gains of more than 40 COMET points for mT5 over the baseline when at least 50% of tokens are noised. Source correction brings smaller performance gains for NLLB and Tower. Source correction has a slight negative effect on GPT-3.5, indicating that extremely robust models have little to gain from source correction. However, Table 12 shows that correcting clean inputs is nearly harmless. mT5 never loses more than COMET points. Even the weaker JamSpell never causes a degradation of more than half a COMET point, suggesting that it could be useful because of its fast runtime.
5.3 Correcting MTNT
Method | OPUS | NLLB | Tower | GPT-3.5 |
Base | 77.51 | 79.20 | 80.65 | 81.33 |
Finetuned | 79.37 | 80.85 | - | - |
JamSpell | 76.42 | 78.09 | 79.17 | 79.68 |
mT5 | 76.41 | 77.88 | 79.34 | 80.16 |
Oracle | 78.65 | 80.11 | 81.59 | 82.16 |
Results.
Social media data is a challenge for both of our source correctors because they were trained on other domains. Using either corrected source leads to a decline of at least one COMET point compared to the raw source. Although an oracle delivers some gains, it still cannot equal the performance of finetuning when it is available. These results seem to confirm that MTNT is not particularly noisy in terms of spelling errors (Karpukhin et al., 2019; Michel and Neubig, 2018).
Per-sentence analysis.
Based on the results in Table 13, it would be easy to conclude that source correction does not work for MTNT. However, Figure 5 shows that some sentences do benefit from correction, and others are not harmed. In these terms, mT5 appears to be a higher-risk corrector than JamSpell, improving translation quality for of examples (versus for JamSpell) while making quality worse for of examples (versus for JamSpell). This, in addition to the oracle performance, suggests that further research could make source correction a useful tool in this domain.
6 Conclusion
We presented several experiments testing the robustness of MT systems to synthetic and natural noise. On synthetic noise, we showed that large multilingual MT models and LLMs are far more robust than older techniques. The experiments on social media translation showed that the same larger models also worked better on natural noise. We added further support for this conclusion through reference-free translation experiments with a novel evaluation metric based on quality estimation. Finally, we exhibited circumstances in which pipeline-based source correction techniques can improve performance on noisy text, both synthetic and natural.
Acknowledgments
This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C645008882- 00000055 (Center for Responsible AI), and by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020.

### v2

@@ 4 Robustness to Social Media Text
The previous experiments show that large translation models and LLMs are more robust to synthetic character perturbations than conventional MT models. But is this result applicable to “authentically noisy” domains such as social media text? The nature of “noise” here is different than in the synthetic task: social media text does not necessarily contain many errors (Rello and Baeza-Yates, 2012), but the domain is very different from FLORES. This makes it difficult to isolate the effect of noise from the general domain adaptation problem. Ideally, we would have a translation corpus in which each example is a triple consisting of an original noisy source, a manually annotated cleaned source, and a gold standard translation. This would allow translations of clean and noisy versions of the same source to be compared on some reference-based metric, isolating the effect of the errors. As we are aware of only one such corpus (Bawden and Sagot, 2023), we instead perform two complementary investigations. First, we evaluate on MTNT (Michel and Neubig, 2018), a noisy social media MT corpus. Although this is a useful test of our models, the noise is not labeled and there is no clean version of the same data to compare to. This motivates our second experiment, in which we translate data from MultiLexNorm (van der Goot et al., 2021), a lexical normalization benchmark. Together, these experiments allow us to see both which models succeed and how badly they fail.
4.1 MTNT Experiments
MTNT pairs Reddit posts with high-quality professional translations. Although the references are somewhat clean, the sources are only lightly filtered, making them potentially noisy.777 percent of MTNT enfr source tokens are misspelled (Michel and Neubig, 2018). This is a higher rate than in formal corpora, but lower than in our synthetic experiments. Unfortunately no cleaned sources exist, making the effect of noise difficult to isolate. Despite this difficulty, it is often used as a robustness benchmark (Karpukhin et al., 2019; Park et al., 2020; Salesky et al., 2021; Vaibhav et al., 2019, inter alia).
Finetuning.
We finetuned OPUS on MTNT enfr as described in Appendix E. We dub this model r/OPUS.
| Method | enfr | fren |
| OPUS | 77.21 | 79.64 |
| r/OPUS | 79.22 | 81.94 |
| NLLB | 79.33 | 80.59 |
| TI | 81.91 | 83.66 |
| GPT-3.5 | 81.33 | 84.72 |
Results.
Results are shown in Table 7. Despite in-domain finetuning benefiting OPUS by more than 2 COMET points for both enfr and fren, this does not close the gap to TI and GPT-3.5. This suggests that TI and GPT-3.5 benefit from their massive training corpora, which likely contain large quantities of social media text. In contrast, the only social media text the finetuned OPUS models have seen are MTNT’s tiny training sets (36k parallel examples for enfr, 19k for fren), plus whatever is in the Tatoeba Challenge corpora.
@@ 4.2 MultiLexNorm Experiments
We use the English, German, and Spanish data from MultiLexNorm as our translation sources. In experiments with English sources, we translate to German and Spanish; otherwise, we translate to English. Statistics are presented in Appendix F.
@@
As MultiLexNorm lacks reference translations, we use three reference-free metrics. Faux-BLEU (Anastasopoulos, 2019) computes (Papineni et al., 2002; NLLB Team et al., 2022), where and are the hypotheses computed from the noisy source and the clean source, respectively. is treated as a pseudoreference. By analogy we also compute faux-COMET. These faux-metrics measure the similarity between and , with faux-BLEU capturing lexical features while faux-COMET is “deeper.” In addition, we use a new metric that we dub QE. Given , , noisy and clean source sequences and , and a reference-free quality estimation metric QE, . We compute QE with COMETKiwi888Specifically, we use Unbabel/wmt22-cometkiwi-da. (Rei et al., 2022b). A QE close to zero means that a model produces similar-quality outputs for both inputs, indicating robustness, whereas a large positive value indicates that translation quality suffers on noisy data.
@@ Results.
Table 8 shows the performance of all models with both noisy and gold-standard cleaned versions of the corpora. In terms of QE, GPT-3.5 performs best for all language pairs. It also records the best faux-BLEU for all pairs except ende, on which TI and NLLB both outperform it. The faux-COMET results show a split, where GPT-3.5 has the highest scores for both to-English pairs but TI passes it for ende and enes.
5 Strategies for Mitigating Noise
So far we have shown that OPUS is less robust to synthetic noise than larger models and performs worse on social media text. Next we evaluate two techniques for mitigating noise: finetuning MT models on synthetically noised data and incorporating a source correction pipeline. These approaches have contrasting trade-offs. Finetuning the MT model allows the robust translation task to be learned end-to-end. However, this is expensive if a model is very large (TI, NLLB) and impossible if its weights are closed (GPT-3.5). On the other hand, pipelines are modular, allowing the same correction system to be reused with any model. The drawback of this modularity is that pipelines may introduce errors as well as fixing them.
5.1 Synthetic Experiments
As a comparison of finetuning and source correction, we focus on enpt with the same synthetically noised corpora as in Section 3. We kept the training data as similar as possible between finetuning and correction experiments: in both cases, we subsampled 5 million examples from the Tatoeba Challenge enpt training set. We noised 15% of source tokens with each of the four noise types used in Section 3, so a total of 60% of tokens were corrupted. For validation, we concatenated the original FLORES enpt development set to a version of the same set in which 20% of tokens have been corrupted with each noise type. By including clean validation data, this favors models that do not forget how to translate clean data. For finetuning OPUS, these noised source examples are paired with clean Portuguese targets. For finetuning correction models, they are paired with the original clean English sources. We followed the training procedure in Appendix E for both MT and correction models.
| Model | clean | swap | drop | dupe | key |
| OPUS | 88.94 | -72.97 | -69.66 | -57.94 | -75.81 |
| +finetuning | 88.52 | -2.14 | -7.59 | -0.87 | -5.01 |
| +correction | 88.36 | -2.02 | -11.81 | -0.08 | -7.20 |
| NLLB | 89.58 | -22.41 | -21.71 | -4.13 | -25.57 |
| +correction | 89.11 | -1.57 | -9.09 | -0.07 | -5.47 |
| TI | 90.02 | -13.44 | -14.68 | -1.89 | -22.04 |
| +correction | 89.57 | -0.86 | -5.59 | -0.16 | -2.99 |
| ChatGPT | 89.83 | -3.76 | -6.63 | -0.98 | -7.78 |
MT Finetuning
Model
clean
swap
drop
dupe
key
100k
88.28
-12.08
-18.96
-4.63
-18.65
1m
88.13
-5.21
-11.72
-2.04
-9.49
5m
88.52
-2.14
-7.59
-0.87
-5.01
Source Correction
Model
clean
swap
drop
dupe
key
100k
88.59
-4.12
-15.31
-0.11
-10.84
1m
88.30
-2.25
-12.83
-0.15
-7.83
5m
88.36
-2.02
-11.81
-0.08
-7.20
Source correction (SC).
As our base model for SC, we adopted ByT5-Small (Xue et al., 2022). At inference time we report the results of a pipeline that pairs the corrector with a translation model. In intrinsic terms, this corrector manages a chrF (Popović, 2015) of at least at correcting each noise type, and over for swaps and dupes. Additional results are shown in Appendix G.
Results.
Our main results are shown in Table 9. For OPUS, both MT finetuning and correction greatly improve robustness to synthetic errors, with both techniques reporting flatter COMET-slopes than GPT-3.5 for all noise types except drop. At a more granular level, in Figure 3 we show OPUS’s COMET scores at all levels of swap noise for MT finetuning and SC. It is clear that MT finetuning outperforms SC across almost all noise levels. Despite NLLB and TI being more robust than OPUS in general, they too benefit from SC, suggesting that its effect is complementary with the models’ inherent robustness. Both MT finetuning and SC slightly degrade model performance on clean translation. However, these degradations are small, and become smaller as the base model gets larger: corrected OPUS declines COMET, versus only for TI. These results also show the surprising difficulty of drop errors: although these are not more problematic for baseline models than other error types (see Tables 2 and 3), neither correction nor MT finetuning handles them as effectively as the other error types. This suggests that the missing information from a single deleted character often cannot be recovered from surrounding context.
Effect of training size.
Although we used 5 million training examples in our main experiments, Table 10 shows the effect that using fewer has on OPUS’s robustness. The correction approach performs much better when data is restricted to 100k examples, but this advantage shrinks as the training size is increased and eventually MT finetuning outperforms it. This supports the intuition that noisy translation is a harder task than correction, so the OPUS model requires more examples to learn it.
How often does mitigation work?
In the main results, we showed that MT finetuning and SC both greatly improve robustness. However, as the scores are presented at corpus level, they do not give insight into what percentage of examples are actually helped, versus how many are harmed. To fill in this missing piece, we compared sentence-level COMET between our baseline models and the two mitigation approaches. Results are shown in Figure 4. At high noise levels, both techniques nearly always help OPUS. However, they behave differently on clean sources. While SC almost never improves clean scores and harms less than 10% of examples, finetuning is more of a high-risk strategy, as it makes scores worse for about half of examples while also improving COMET for 38% of them. The trend is subtly different for SC with TI. While correction helps more examples than it hurts whenever the noise level is at least 10%, the percentage of examples that are harmed by correction actually increases up to 30% noise. This suggests that correction may introduce errors into some noisy sequences that TI could have handled itself.
5.2 Mitigating Errors in MTNT
Having shown that MT finetuning and SC are both effective techniques for improving robustness to synthetic errors, we return to MTNT. Our goal is to determine whether either MT finetuning or SC can provide some benefit to performance on enfr.
| Finetuning | Uncorrected | Corrected |
| None | 77.21 | 76.31 |
| MTNT | 79.22 | 77.97 |
| Synthetic | 75.95 | 75.31 |
Finetuning vs. correction for OPUS.
We evaluate MT finetuning and SC on MTNT. For MT finetuning, we compare two techniques: finetuning only on MTNT, as in Section 4.1; and finetuning on 5 million synthetically noised Tatoeba Challenge enfr examples, matching the procedure used for enpt in Section 5.1. For SC, we use the same English model as in the synthetic experiments. Results are shown in Table 11. Finetuning on MTNT outperforms the baseline, but both SC and MT finetuning are harmful, confirming that MTNT is not very noisy in terms of spelling errors (Karpukhin et al., 2019; Michel and Neubig, 2018).
Effect of source correction.
In Table 12, we show results with and without SC for various models. Although correction does not improve results on the whole, the oracle consistently outperforms the baseline by about COMET, illustrating that many sequences do benefit from correction. Indeed, 21.8% of MTNT examples improve with correction and another 46.1% are not harmed. Future work could investigate the use of a routing mechanism that decides for each example whether to apply SC.
| Method | r/OPUS | NLLB | TI |
| Base | 79.22 | 79.33 | 81.91 |
| Correction | 77.97 | 78.32 | 80.94 |
| Oracle | 79.71 | 79.82 | 82.43 |
6 Conclusion
We presented several experiments testing the robustness of MT systems to synthetic and natural noise. On synthetic noise, we showed that large multilingual MT models and LLMs are far more robust than older models. The experiments on social media translation showed that larger models also worked better on natural noise. We further supported this conclusion through reference-free translation experiments with a novel evaluation metric based on quality estimation. Finally, we showed that noisy finetuning and source correction allow smaller models to exceed GPT-3.5’s robustness synthetic noise, while also being useful in many cases for handling natural errors.
Limitations
We acknowledge the limitations of our work. All of the languages included in this study have large speaker populations and many resources available, and are the official languages of countries. Some conclusions may not generalize to low resource languages. Additionally, this paper studies only one source of natural noise, namely social media text. Other varieties of text perceived as noisy, such as transcribed speech or text written by nonfluent language users, may have different properties. Finally, the evaluation techniques used in this work are all automatic or neural, and may differ from gold-standard human evaluation.
Acknowledgments
This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações. We thank Sweta Agrawal and the anonymous reviewers for their helpful comments.
Appendix A Key Noise
As different languages customarily use different keyboard layouts, we made slight alterations to our key noising procedure for each source language. We use the QWERTZ layout for German, AZERTY for French, QWERTY for English and Portuguese, and South Korean Dubeolsik for Korean. For Korean, we used hangul-jamo999https://github.com/jonghwanhyeon/hangul-jamo to decompose hangul characters into jamo, which represent individual keystrokes, before applying perturbations.
Appendix B OPUS Models
| LP | Vocab | Params |
| ende | 58.1k | 74.4M |
| deen | 58.1k | 74.4M |
| enes | 55.0k | 234.8M |
| esen | 65.0k | 77.9M |
| enfr | 53.0k | 232.7M |
| fren | 53.0k | 232.7M |
| enko | 32.0k* | 209.2M |
| koen | 32.0k* | 209.2M |
| enpt | 54.7k | 234.5M |
| pten | 60.0k | 75.4M |
The parameters and vocabulary sizes of the OPUS models are shown in Table 13. All checkpoints come from the Tatoeba Challenge (Tiedemann, 2020).101010https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/models
Appendix C Inference
Owing to the differing frameworks between OPUS, OPUSLLM, NLLB, TI, and ByT5 models, we use different beam search implementations depending on the model. For OPUS models, we decode with Marian (Junczys-Dowmunt et al., 2018), while NLLB and ByT5 use Hugging Face transformers.111111https://github.com/huggingface/transformers TI models and OPUSLLM are decoded with vllm (Kwon et al., 2023).121212https://github.com/vllm-project/vllm Regardless of framework, we used a beam size of across all experiments.
C.1 Prompt for TowerInstruct
‘‘Translate the following text from [source language] to [target language].nSource:[source text]n[target language]:’’
C.2 Prompt for GPT-3.5
‘‘Translate this sentence from [source language] to [target language].nSource:[source text]nTarget:’’
Appendix D OPUSLLM Training Details
The OPUSLLM model was trained with used Megatron-DeepSpeed131313https://github.com/microsoft/Megatron-DeepSpeed for a maximum of 300k steps with an effective batch size of 65k tokens and a base learning rate of , with a constant learning rate schedule and 5000 warmup steps. The model used the same 32k tokenizer as TI. To account for the model’s decoder-only structure, examples were formatted with two special tokens to indicate the beginning of the source and target sequences. At inference time, we selected the checkpoint with the best COMET on the FLORES enpt dev set.
Appendix E Training Hyperparameters
In this work, we finetune two types of base models: OPUS models and ByT5-Small. Despite the various technical differences between these two models, we used almost identical training procedures for them, with differences noted below.
Basic training procedure.
We finetuned using early stopping with patience 3. We validated every 500 steps. We used a grid over the learning rates and selected the best checkpoint by validation loss.
Training library.
For OPUS models, we finetuned with the marian command line tool from Marian (Junczys-Dowmunt et al., 2018). For ByT5-Small, we used a script that leverages the Trainer class from Hugging Face transformers.
Appendix F MultiLexNorm Statistics
Appendix G Intrinsic Correction Performance
| swap | drop | dupe | key | |
@@ Appendix G Intrinsic Correction Performance
| ByT5-Small | 97.5 | 89.6 | 99.6 | 94.4 |
Intrinsic correction results in terms of chrF (Popović, 2015) are shown in Table 15. In addition to our ByT5-Small corrector, we include the results of the pretrained English model from JamSpell141414https://github.com/bakwc/JamSpell (Ozinov, 2019), a trigram-based spell-checker.
Appendix H Synthetic Results with Other Metrics
The models’ synthetic performance on clean data is shown in terms of BLEU151515nrefs:1|case:mixed|eff:no|tok:flores200|
smooth:exp|version:2.5.1 in Table 16 and in chrF161616nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|
version:2.5.1 in Table 17.
Synthetic results showing BLEU-slope are shown in Tables 18 and 19.
Synthetic results showing chrF-slope are shown in Tables 20 and 21.
BLEU-slope and chrF-slope are computed analogously to COMET-slope.
xxen
Model
deen
fren
koen
pten
avg.
OPUS
44.7
49.7
30.7
49.3
43.6
NLLB
49.3
50.3
34.8
55.1
47.4
TI
50.6
52.1
37.0
56.7
49.1
GPT-3.5
48.8
49.1
32.5
53.6
46.0
enxx
Model
ende
enfr
enko
enpt
avg.
OPUS
42.0
55.6
23.8
54.6
44.0
NLLB
46.8
56.2
24.8
54.5
45.6
TI
47.0
57.1
29.5
52.8
46.6
GPT-3.5
47.8
56.6
25.6
56.6
46.7
xxen
Model
deen
fren
koen
pten
avg.
OPUS
66.4
69.1
55.8
69.4
65.2
NLLB
69.2
69.9
58.6
72.9
67.7
TI
69.9
70.6
60.3
73.9
68.7
GPT-3.5
69.8
69.7
58.2
72.7
67.6
enxx
Model
ende
enfr
enko
enpt
avg.
OPUS
62.9
71.7
36.2
71.3
60.5
NLLB
65.7
71.8
36.7
71.0
61.3
TI
66.2
72.3
38.7
70.3
61.9
GPT-3.5
67.1
72.6
35.1
72.8
61.9
swap
Model
deen
fren
koen
pten
avg.
OPUS
-51.4
-58.2
-28.2
-56.2
-48.5
NLLB
-33.4
-35.8
-30.6
-37.5
-34.3
TI
-37.9
-38.8
-30.4
-39.6
-36.7
GPT-3.5
-10.9
-14.4
-27.4
-17.2
-17.5
drop
Model
deen
fren
koen
pten
avg.
OPUS
-45.6
-49.6
-26.4
-49.1
-42.7
NLLB
-29.8
-29.9
-26.1
-32.7
-29.6
TI
-29.5
-29.7
-25.6
-30.0
-28.7
GPT-3.5
-14.4
-13.8
-24.9
-16.0
-17.3
dupe
Model
deen
fren
koen
pten
avg.
OPUS
-34.3
-38.0
-8.8
-37.6
-29.7
NLLB
-13.9
-15.7
-11.5
-14.8
-14.0
TI
-9.9
-13.0
-5.5
-10.5
-9.7
GPT-3.5
-5.5
-5.3
-12.0
-5.8
-7.2
key
Model
deen
fren
koen
pten
avg.
OPUS
-50.3
-57.3
-27.8
-57.2
-48.2
NLLB
-35.2
-36.4
-33.3
-40.3
-36.3
TI
-38.5
-42.7
-27.2
-46.3
-38.7
GPT-3.5
-18.3
-18.2
-22.5
-21.9
-20.2
swap
Model
ende
enfr
enko
enpt
avg.
OPUS
-49.1
-62.6
-28.9
-64.5
-51.3
NLLB
-31.2
-33.8
-18.1
-37.1
-30.0
TI
-19.8
-20.7
-12.6
-20.0
-18.3
GPT-3.5
-7.8
-9.3
-7.1
-5.1
-7.3
drop
Model
ende
enfr
enko
enpt
avg.
OPUS
-47.8
-59.4
-27.9
-61.4
-49.1
NLLB
-30.7
-33.7
-16.6
-36.2
-29.3
TI
-21.9
-23.5
-13.7
-22.6
-20.4
GPT-3.5
-12.0
-13.4
-9.6
-12.8
-12.0
dupe
Model
ende
enfr
enko
enpt
avg.
OPUS
-39.5
-45.9
-26.5
-54.5
-41.6
NLLB
-9.3
-8.9
-3.6
-14.7
-9.1
TI
-4.6
-5.0
-3.5
-5.2
-4.6
GPT-3.5
-3.7
-4.0
-3.2
-3.1
-3.5
key
Model
ende
enfr
enko
enpt
avg.
OPUS
-48.8
-62.9
-29.0
-64.6
-51.3
NLLB
-34.1
-37.4
-20.1
-40.3
-33.0
TI
-25.0
-26.9
-15.3
-26.8
-23.5
GPT-3.5
-11.8
-13.1
-10.5
-12.5
-12.0
swap
Model
deen
fren
koen
pten
avg.
OPUS
-45.2
-47.4
-32.9
-46.0
-42.9
NLLB
-25.3
-26.6
-27.7
-26.4
-26.5
TI
-28.4
-27.3
-29.0
-28.1
-28.2
GPT-3.5
-7.1
-9.6
-29.2
-10.9
-14.2
drop
Model
deen
fren
koen
pten
avg.
OPUS
-40.1
-39.7
-29.0
-38.7
-36.9
NLLB
-23.8
-23.6
-26.6
-24.1
-24.5
TI
-22.3
-20.9
-25.5
-21.1
-22.5
GPT-3.5
-10.0
-9.3
-26.0
-10.4
-13.9
dupe
Model
deen
fren
koen
pten
avg.
OPUS
-22.9
-24.5
-6.8
-23.1
-19.3
NLLB
-7.6
-8.6
-6.1
-7.5
-7.4
TI
-5.1
-6.4
-3.9
-5.3
-5.2
GPT-3.5
-2.7
-2.7
-10.2
-2.8
-4.6
key
Model
deen
fren
koen
pten
avg.
OPUS
-43.6
-45.5
-34.5
-46.0
-42.4
NLLB
-26.6
-27.7
-31.0
-29.6
-28.7
TI
-28.6
-30.6
-26.3
-34.6
-30.0
GPT-3.5
-12.1
-11.9
-22.6
-14.0
-15.2
swap
Model
ende
enfr
enko
enpt
avg.
OPUS
-50.0
-56.8
-38.0
-55.7
-50.1
NLLB
-24.7
-26.4
-21.0
-27.4
-24.9
TI
-15.9
-15.0
-11.8
-14.8
-14.4
GPT-3.5
-5.1
-6.0
-6.4
-5.1
-5.7
drop
Model
ende
enfr
enko
enpt
avg.
OPUS
-48.3
-53.3
-36.6
-53.5
-47.9
NLLB
-24.7
-26.6
-20.3
-27.2
-24.7
TI
-17.4
-17.0
-12.7
-16.6
-15.9
GPT-3.5
-8.3
-9.0
-8.4
-8.2
-8.5
dupe
Model
ende
enfr
enko
enpt
avg.
OPUS
-34.7
-36.8
-33.6
-41.8
-36.7
NLLB
-5.7
-5.8
-5.7
-8.6
-6.4
TI
-2.8
-2.6
-2.8
-2.7
-2.7
GPT-3.5
-1.9
-2.1
-2.5
-1.3
-2.0
key
Model
ende
enfr
enko
enpt
avg.
OPUS
-50.0
-56.6
-37.9
-55.9
-50.1
NLLB
-27.9
-29.8
-22.7
-30.4
-27.7
TI
-21.0
-20.6
-15.3
-20.5
-19.4
GPT-3.5
-8.4
-9.0
-8.8
-8.1
-8.6